OpenStudy (anonymous):

What is the slope of the equation y =5/4x-7/4

OpenStudy (kidrah69):

y=mx+b m=slope

OpenStudy (kidrah69):

Do you know what it is now?

OpenStudy (anonymous):

m=5/4 m=slope

OpenStudy (kidrah69):

yep :)

OpenStudy (kidrah69):

Welcome to openstudy!